# RAMEN: Real-time Asynchronous Multi-agent Neural Implicit Mapping

### v1

@@ Abstract
Multi-agent neural implicit mapping allows robots to collaboratively capture and reconstruct complex environments with high fidelity. However, existing approaches often rely on synchronous communication, which is impractical in real-world scenarios with limited bandwidth and potential communication interruptions. This paper introduces RAMEN: Real-time Asynchronous Multi-agEnt Neural implicit mapping, a novel approach designed to address this challenge. RAMEN employs an uncertainty-weighted multi-agent consensus optimization algorithm that accounts for communication disruptions. When communication is lost between a pair of agents, each agent retains only an outdated copy of its neighbor’s map, with the uncertainty of this copy increasing over time since the last communication. Using gradient update information, we quantify the uncertainty associated with each parameter of the neural network map. Neural network maps from different agents are brought to consensus on the basis of their levels of uncertainty, with consensus biased towards network parameters with lower uncertainty. To achieve this, we derive a weighted variant of the decentralized consensus alternating direction method of multipliers (C-ADMM) algorithm, facilitating robust collaboration among agents with varying communication and update frequencies. Through extensive evaluations on real-world datasets and robot hardware experiments, we demonstrate RAMEN’s superior mapping performance under challenging communication conditions.
@@ III-A Single-agent Neural Implicit Mapping
We train the neural implicit map by minimizing the difference between rendered RGB-D images and the images from robot’s camera. We use notation $R$ to denote the actual images collected by the robot. The objective function is then given as:
@@ III-B Synchronous Multi-agent Mapping
We now discuss how Co-SLAM can be extended into synchronous multi-agent mapping. In a multi-agent setup, a group of agents collaboratively map their environment and share information within a communication graph $G=(\mathcal{V},\mathcal{E})$. Each agent is represented as a node $i\in\mathcal{V}$, and their communication connectivity is defined by a set of edges $\mathcal{E}$. Each agent $i$ can communicate with its neighbors $N_{i}=\{j\in\mathcal{V}\mid(i,j)\in\mathcal{E}\}$, sharing its learned parameters $\mathbf{\Theta}_{i}$ instead of raw images. Each agent $i$ captures posed RGB-D images $R_{i}$ with its onboard camera. During each mapping iteration, agent $i$ samples pixels from $R_{i}$ and minimizes its local objective function $L^{obj}_{i}$:
@@ IV Asynchronous Multi-agent Neural Implicit Mapping
This section details the core components of RAMEN’s asynchronous multi-agent mapping framework. We first derive an uncertainty-weighted version C-ADMM algorithm, utilizing uncertainty to favor reliable information in the consensus optimization process. Then, we introduce our proposed method for capturing uncertainty is a frequency-based epistemic uncertainty quantification method, detailing how it assesses the reliability of each agent’s contributions to the collaborative mapping process.
@@ IV-A Uncertainty-Weighted Decentralized C-ADMM
where we use a weighted norm $\|\mathbf{x}\|^{2}_{\mathbf{W}}=\mathbf{x}^{\top}\mathbf{W}\mathbf{x}$ instead of L2 norm in (8). Note that $\mathbf{W}_{ij}^{t},\mathbf{W}_{ji}^{t}$ represent weights reflecting the relative reliability of information shared between agents i and j at iteration t. Intuitively, a higher weight indicates greater confidence/lower uncertainty in the corresponding information
Applying ADMM [42] to (10), at each iteration $t$ for agent $i$, we alternatively update local consensus variable $\mathbf{z}_{ij}$, neural implicit map $\mathbf{\Theta}_{i}$, and consensus-enforcing dual variables $\boldsymbol{\lambda}_{ij},\mathbf{v}_{ij}$. We obtain $\mathbf{z}_{ij}^{t+1}$ by setting the derivative of (10) with respect to $\mathbf{z}_{ij}$ to zero to get:
@@ IV-B Frequency-based Epistemic Uncertainty
Based on these two observations, we associate the number of gradient updates a feature grid parameter receives with its epistemic uncertainty. More updates indicate more frequent visits to the corresponding location, implying lower uncertainty. Due to the correlated nature of parameters in MLP decoders $F_{\tau},F_{\phi}$ and their lack of direct spatial correspondence, we pretrain the decoders using the NICE-SLAM apartment dataset [23] and keep their parameters fixed during mapping. Thus, unlike Co-SLAM, we only optimize the parameters $\alpha$ in the feature grid $V_{\alpha}$, resulting in $\mathbf{\Theta}=\{\alpha\}$.
@@
$\mathbf{u}_{i}^{T}=\sum_{t=0}^{T}\text{sgn}\left(\text{abs}\left(\frac{\partial L^{obj}_{i,t}}{\partial\mathbf{\Theta}_{i}^{t}}\right)\right),$ (18)
where $\mathbf{u}_{i}^{T}$ is the feature grid uncertainty of agent $i$ that is accumulated until time $T$; $L^{obj}_{i,t}$ is the Co-SLAM reconstruction objective loss (6) of agent $i$ at iteration $t$; $\text{abs}(\cdot)$ is the absolute value function; and $\text{sgn}(\cdot)$ is the sign function returning 1 when the given value $>0$ and returning 0 otherwise. Our uncertainty quantification (18) links the uncertainty of each feature grid parameter to its corresponding accumulated number of non-zero gradient updates. Fig 2 visualizes the uncertainty of the parameters in the coarsest feature grid $V_{\alpha}^{l=1}$ in an example using our proposed method, demonstrating that our method accurately captures the uncertainty of the neural implicit map.
Since the number of gradient updates could be infinite, we cannot directly use $\mathbf{u}_{i}^{T}$ as weights for consensus optimization in IV-A. Instead, we normalize the uncertainty values to obtain bounded weights. Given $\mathbf{u}_{i}^{T_{i}}$ and $\mathbf{u}_{j}^{T_{j}}$ for agent $i$ and its neighbor $j$, we normalize them into weights $\mathbf{W}_{ij}$ and $\mathbf{W}_{ji}$ as follows:
$\mathbf{u}_{sum}:=\mathbf{u}_{i}^{T_{i}}+\mathbf{u}_{j}^{T_{j}},$ (19a)
$\zeta=\beta_{l}-\epsilon\cdot\text{min}(\mathbf{u}_{sum}),$ (19c)
$\mathbf{W}_{ij}=\epsilon\cdot\mathbf{u}_{i}^{T_{i}}+\zeta,\mathbf{W}_{ji}=\epsilon\cdot\mathbf{u}_{j}^{T_{j}}+\zeta,$ (19d)
where $\beta_{u}$ and $\beta_{l}$ are the user-specified upper and lower bounds of the weights, $\text{max}(\cdot)$ and $\text{min}(\cdot)$ return the maximum and the minimum elements of the input vector, and $\epsilon$ (19b) and $\zeta$ (19c) are scaling and bias terms for normalization. The normalization (19) preserves the proportional relationship between the uncertainty values of agents $i$ and $j$ by applying the same $\epsilon$ and $\zeta$ to both $\mathbf{u}_{i}^{T_{i}}$ and $\mathbf{u}_{j}^{T_{j}}$. These weights are then incorporated into the weighted decentralized C-ADMM algorithm to effectively leverage information from neighboring models while accounting for their uncertainty.
@@ V-B Real-World Hardware Experiments
To evaluate RAMEN’s performance in real-time multi-robot mapping, we used two Turtlebots within the Robot Operating System (ROS) [52]. Each robot was manually controlled to explore half of the environment, capturing RGBD images with its onboard sensor. Figure 1 illustrates the trajectories and areas covered by each robot. In this setup, due to the limited range of the stereo camera, successful consensus between the robots is crucial for reconstructing the complete scene. We used a VICON motion capture system to provide ground-truth robot poses. Images and pose data from each Turtlebot were streamed to a separate PC via ROS2 and WiFi for real-time processing, with mapping optimization performed at 1Hz. Due to bandwidth limitations, the two PCs could only exchange their neural implicit maps every 30 seconds, resulting in a communication success rate of 3.33%. This setup, combined with noisy depth images, motion blur, and synchronization errors between images and pose data, presents an extremely challenging scenario that thoroughly tests RAMEN’s robustness.
@@ VI Limitations
A key limitation of the proposed frequency-based epistemic uncertainty is its inability to differentiate between regions with varying geometric complexity. Regions with complex geometries, such as those containing many small objects or fine details, require more observations to reduce uncertainty and improve reconstruction quality. Another limitation, inherited from Co-SLAM, is the reliance on a keyframe database $R_{i}$. This necessitates storing a large number of images onboard and continuously replaying them to prevent catastrophic forgetting, posing challenges for large-scale mapping tasks. Finally, controlling the robots manually in our hardware experiments may have introduced a human bias, potentially affecting reproducibility. A promising future direction is to integrate active neural implicit mapping techniques, such as those proposed in [36], to automate data acquisition.

### v2

@@ Abstract
Multi-agent neural implicit mapping allows robots to collaboratively capture and reconstruct complex environments with high fidelity. However, existing approaches often rely on synchronous communication, which is impractical in real-world scenarios with limited bandwidth and potential communication interruptions. This paper introduces RAMEN: Real-time Asynchronous Multi-agEnt Neural implicit mapping, a novel approach designed to address this challenge. RAMEN employs an uncertainty-weighted multi-agent consensus optimization algorithm that accounts for communication disruptions. When communication is lost between a pair of agents, each agent retains only an outdated copy of its neighbor’s map, with the uncertainty of this copy increasing over time since the last communication. Using gradient update information, we quantify the uncertainty associated with each parameter of the neural network map. Neural network maps from different agents are brought to consensus on the basis of their levels of uncertainty, with consensus biased towards network parameters with lower uncertainty. To achieve this, we derive a weighted variant of the decentralized consensus alternating direction method of multipliers (C-ADMM) algorithm, facilitating robust collaboration among agents with varying communication and update frequencies. Through extensive evaluations on real-world datasets and robot hardware experiments, we demonstrate RAMEN’s superior mapping performance under challenging communication conditions. Supplementary videos and codes are available at https://iconlab.negarmehr.com/RAMEN/.
@@ III-A Single-agent Neural Implicit Mapping
We train the neural implicit map by minimizing the difference between rendered RGBD images and the images from robot’s camera. We use notation $R$ to denote the actual images collected by the robot. The objective function is then given as:
@@ III-B Synchronous Multi-agent Mapping
We now discuss how Co-SLAM can be extended into synchronous multi-agent mapping. In a multi-agent setup, a group of agents collaboratively map their environment and share information within a communication graph $G=(\mathcal{V},\mathcal{E})$. Each agent is represented as a node $i\in\mathcal{V}$, and their communication connectivity is defined by a set of edges $\mathcal{E}$. Each agent $i$ can communicate with its neighbors $N_{i}=\{j\in\mathcal{V}\mid(i,j)\in\mathcal{E}\}$, sharing its learned parameters $\mathbf{\Theta}_{i}$ instead of raw images. Each agent $i$ captures posed RGBD images $R_{i}$ with its onboard camera. During each mapping iteration, agent $i$ samples pixels from $R_{i}$ and minimizes its local objective function $L^{obj}_{i}$:
@@ IV Asynchronous Multi-agent Neural Implicit Mapping
This section details the core components of RAMEN’s asynchronous multi-agent mapping framework. In our context, asynchronism arises from two primary sources: communication drops and variations in processing speed. This work primarily addresses communication drops, as they pose a significant challenge for both homogeneous and heterogeneous robot teams. We first derive an uncertainty-weighted version C-ADMM algorithm, utilizing uncertainty to favor reliable information in the consensus optimization process. Then, we introduce our proposed frequency-based epistemic uncertainty quantification method, detailing how it assesses the reliability of each agent’s contributions to the collaborative mapping process.
@@ IV-A Uncertainty-Weighted Decentralized C-ADMM
where we use a weighted norm $\|\mathbf{x}\|^{2}_{\mathbf{W}}=\mathbf{x}^{\top}\mathbf{W}\mathbf{x}$ instead of L2 norm in (8). Note that $\mathbf{W}_{ij}^{t},\mathbf{W}_{ji}^{t}$ are weight matrices reflecting the relative reliability of information shared between agents i and j at iteration t. Intuitively, a higher weight indicates greater confidence/lower uncertainty in the corresponding information.
Applying ADMM [42] to (10), at each iteration $t$ for agent $i$, we alternatively update local consensus variable $\mathbf{z}_{ij}$, neural implicit map $\mathbf{\Theta}_{i}$, and consensus-enforcing dual variables $\boldsymbol{\lambda}_{ij},\mathbf{v}_{ij}$. We obtain $\mathbf{z}_{ij}^{t+1}$ by setting the derivative of (10) with respect to $\mathbf{z}_{ij}$ to zero:
@@ IV-B Frequency-based Epistemic Uncertainty
Based on these two observations, we associate the number of gradient updates a feature grid parameter receives with its epistemic uncertainty. More updates indicate more frequent visits to the corresponding location, implying lower uncertainty. Due to the correlated nature of parameters in MLP decoders $F_{\tau},F_{\phi}$ and their lack of direct spatial correspondence, we pretrain the decoders using the NICE-SLAM apartment dataset [23] and keep their parameters fixed during mapping. Thus, unlike Co-SLAM, we only optimize the parameters $\alpha$ in the feature grid $V_{\alpha}$, resulting in $\mathbf{\Theta}=\{\alpha\}$ with $n$ learnable parameters. We flatten $\mathbf{\Theta}$ into a $n\times 1$ vector in our computations.
@@
$\mathbf{u}_{i}^{t}=\sum_{k=0}^{t}\text{sgn}\left(\text{abs}\left(\frac{\partial L^{obj}_{i,k}}{\partial\mathbf{\Theta}_{i}^{k}}\right)\right),$ (18)
where $\mathbf{u}_{i}^{t}$, a $n\times 1$ vector, represents the uncertainty associated with each feature grid parameter at iteration $t$. $L^{obj}_{i,k}$ is the Co-SLAM reconstruction objective loss (6) of agent $i$ at iteartion $k$; $\text{abs}(\cdot)$ is the absolute value function; $\text{sgn}(\cdot)$ is the sign function returning 1 when the given value $>0$ and returning 0 otherwise. $\mathbf{u}_{i}^{t}$ tracks the count of non-zero gradient updates for each parameter, establishing a link between the frequency of spatial location visits and their uncertainty. Fig 2 visualizes the uncertainty of the parameters in the coarsest feature grid $V_{\alpha}^{l=1}$ in an example using our proposed method, demonstrating that our method accurately captures the uncertainty of the neural implicit map.
@@
Since the number of gradient updates could be infinite, we cannot directly use $\mathbf{u}_{i}^{t}$ as weights for consensus optimization in IV-A. Simply normalizing $\mathbf{u}_{i}^{t_{i}}$ and $\mathbf{u}_{j}^{t_{j}}$ independently to a range of $[0,1]$ would eliminate the proportional relationship between their elements. For example, the maximum value in $\mathbf{u}_{i}^{t_{i}}$ might be substantially larger than the maximum in $\mathbf{u}_{j}^{t_{j}}$, a difference that would be lost after independent normalization. To preserve these proportional relationships and ensure that the weights reflect the relative uncertainties, we employ a linear scaling and shifting operation. Given $\mathbf{u}_{i}^{t_{i}}$ and $\mathbf{u}_{j}^{t_{j}}$ for agent $i$ and its neighbor $j$, we obtain $n\times n$ weight matrices $\mathbf{W}_{ij}$ and $\mathbf{W}_{ji}$ as follows:
$\mathbf{u}_{sum}:=\mathbf{u}_{i}^{t_{i}}+\mathbf{u}_{j}^{t_{j}},$ (19a)
$\zeta=\beta_{l}-\epsilon\cdot\text{min}(\mathbf{u}_{sum}),$ (19c)
$\mathbf{W}_{ij}=\text{diag}(\epsilon\cdot\mathbf{u}_{i}^{t_{i}}+\zeta),\mathbf{W}_{ji}=\text{diag}(\epsilon\cdot\mathbf{u}_{j}^{t_{j}}+\zeta),$ (19d)
where $\beta_{u}$ and $\beta_{l}$ are the user-specified upper and lower bounds of the weights; $\text{max}(\cdot)$ and $\text{min}(\cdot)$ return the maximum and the minimum elements of the input vector; $\text{diag}(\cdot)$ returns a $n\times n$ matrix using the input as its diagonal elements; $\epsilon$ (19b) and $\zeta$ (19c) are scaling and shifting terms for normalization. The normalization (19) preserves the proportional relationship between the uncertainty values of agents $i$ and $j$ by applying the same $\epsilon$ and $\zeta$ to both $\mathbf{u}_{i}^{t_{i}}$ and $\mathbf{u}_{j}^{t_{j}}$. These weight matrices are then incorporated into the weighted decentralized C-ADMM algorithm to effectively leverage information from neighboring models while accounting for their uncertainty.
@@ V-A Simulated Experiments
We also emphasize that RAMEN employs a simple uncertainty quantifier, incurring minimal computation overhead compared to existing NeRF uncertainty methods covered in Section II. For instance, in a representative three-agent experiment conducted on an RTX 4090 GPU, RAMEN runs at 8.58 FPS, while the baseline DiNNO runs at 9.38 FPS.
@@ V-B Real-World Hardware Experiments
To evaluate RAMEN’s performance in real-time multi-robot mapping, we used two Turtlebots within the Robot Operating System (ROS) [52]. Each robot was manually controlled to explore half of the environment, capturing RGBD images with its onboard sensor. Figure 1 illustrates the trajectories and areas covered by each robot. In this setup, due to the limited range of the stereo camera, successful consensus between the robots is crucial for reconstructing the complete scene. We used a VICON motion capture system to provide ground-truth robot poses. Images and pose data from each Turtlebot were streamed to a separate PC via ROS2 and WiFi for real-time processing, with mapping optimization performed at around 1Hz. Two PCs exhibited substantial differences in processing speed, and no synchronization mechanism was implemented to align their optimization iteration counts. Due to bandwidth limitations, the two PCs could only exchange their neural implicit maps every 30 seconds, resulting in a communication success rate of 3.33%. This asynchronous communication setup, combined with noisy depth images and motion blur, presents an extremely challenging scenario that thoroughly tests RAMEN’s robustness.
@@ VI Limitations
A key limitation of the proposed frequency-based epistemic uncertainty is its inability to differentiate between regions with varying geometric complexity. Regions with complex geometries, such as those containing many small objects or fine details, require more observations to reduce uncertainty and improve reconstruction quality. Another limitation, inherited from Co-SLAM, is the reliance on a keyframe database $R_{i}$. This necessitates storing a large number of images onboard and continuously replaying them to prevent catastrophic forgetting, posing challenges for large-scale mapping tasks. Our reconstruction quality also suffer from significant depth estimation noise introduced by the RGBD camera on the Turtlebot. Finally, controlling the robots manually in our hardware experiments may have introduced a human bias, potentially affecting reproducibility. A promising future direction is to integrate active neural implicit mapping techniques, such as those proposed in [36], to automate data acquisition.
@@ VII Conclusion
Acknowledgments
This work is supported in part by the National Science Foundation, under grants ECCS-2438314 CAREER Award, CNS-2423130, and CCF-2423131.